OpenStudy (anonymous):

Explain the process of simplifying the following expression and the final simplified expression. 4 *square root* 12a^4b^2

OpenStudy (anonymous):

\[4\sqrt{12a^{4}b ^{2}}\] like this ? or \[\sqrt[4]{12a ^{4}b ^{2}}\] this?

jhonyy9 (jhonyy9):

i think the first one wann being @mebs

OpenStudy (anonymous):

So remember that 4 x 3 = 12 so we can write \[\sqrt{12} = 2\sqrt{3}\] the a term is written as \[a ^{4-2} = a ^{2}\] The b term is written as b. We get the expression \[8\sqrt{3} a ^{2}b\]

jhonyy9 (jhonyy9):

@mebs for not being possible confusing write hence 8a^2bsqrt(3)

OpenStudy (anonymous):

Thanks your right should answer in the same format.@jhonyy9